### v1

@@ 4.3 Spectral Energy Distribution
We queried the SED of WeSb 1 using the online VOSA SED tool (Bayo et al., 2008) (Figure 10, blue squares) with a 3′′ cross-match radius with all the available catalogs. The flux outputs are presented in Table 4. VOSA failed to load the GALEX near-ultra-violet (NUV) flux, which was separately queried from the GALEX catalog161616https://galex.stsci.edu/GR6/?page=mastform and manually added to the table.
First, we analyze the SED independent of the inferences about the stellar parameters/spectral type in the previous section. We consider three cases of extinction: only interstellar (171717Obtained from the online query service: http://argonaut.skymaps.info/query, implementing the Green et al. (2019) three-dimensional extinction maps), a representative value for the Balmer decrement (), and the Na-D absorption (, as an example of an extreme case of extinction).181818We note here that the EW of the Na-D absorption in the Coelho spectrum is . Subtracting this from the observed EW leads to a revised extinction estimation of , consistent with all the other estimates. We still briefly consider this case just to demonstrate the effect of a large extinction, in case it is underestimated. The first three panels of Figure 10 show the reddening-corrected SEDs for these three extinction values. A strong IR excess, starting from the near-infrared wavelengths is evident, likely arising from circumstellar dust disk and/or shell. The stellar component dominates in the optical wavelengths.
@@ 4.4 Discussion
The position of the star in the Gaia color-magnitude diagram (CMD), however, is not consistent with a single white dwarf. With Gaia and , the object is significantly redder and more luminous than a white dwarf. All the previous analyses substantiate this, making it more likely to be a binary with the optical flux dominated by a cooler companion (A – G type star) of the hot CSPN. Thus, the observed photometric variability can arise from a transiting circumbinary debris or a circumstellar debris disk around the companion. The near-total eclipses, however, demands that the obscuring material be nearly as large as the star itself. With the inferred stellar radius being , it is more likely that the transits are from very thick dust/debris clouds rather than solid objects like asteroids. We note here that dust activity from the star itself might also have resulted in irregular variability (with RCrB stars being extreme cases191919Note that the WeSb 1 nucleus is unlikely to be an RCrB or similar hydrogen deficient star due to the presence of prominent Balmer/Paschen absorption lines). However, the best estimate of the spectral type of the companion, F-type subgiant, makes this scenario unlikely as such stars are not expected to take part in significant dust production.
@@
The inferred temperature and radius of the companion can be used to estimate its mass and the age of the system. We query stellar evolutionary tracks calculated using the Modules for Experiments in Stellar Astrophysics (MESA, Paxton et al. 2011, 2013, 2015) from the online portal of MESA Isochrones and Stellar Tracks (MIST, Dotter 2016; Choi et al. 2016)202020https://waps.cfa.harvard.edu/MIST/index.html. We visually compare the models with the stellar parameters calculated in the previous sections. The result is presented in Figure 11. We find that the inferred radius and temperature values lie between stars of masses 2.2 – 2.4 (left panel in the figure). With a generous allowance for error, we arrive at a mass estimation of . Thus, we get a corresponding age of the star of Gyr (middle panel in the figure). Note that at this evolutionary stage, the stellar radius (and temperature) evolves rapidly with time. Thus, the age estimation is not particularly sensitive to the exact estimations of the stellar parameters. At this mass and age, the model predicts , consistent with the previous estimates. The position of the star in the Gaia CMD is also consistent with those predicted by the models in the inferred mass range (with an applied extinction correction of , see right panel of the same figure). As the PN is still visible, this implies that the PN progenitor left the main sequence only recently. Thus, its initial mass needs to be very similar (slightly larger) to its subgiant companion. We thus obtain estimates of both the total initial mass of the binary and its age. We also find that the exemplary stellar parameters of the hot CSPN used in Figure 10 are also broadly consistent with this evolutionary stage (see Appendix F). We note here, however, that there is a chance that the companion is not naturally in the subgiant phase (as indicated by the radius), but is just inflated from capture of the nebular ejecta. In that case, the mass and age estimation will no longer be valid.
@@
The compactness of the emission lines and the high inferred electron densities (see section 4.2) are additional puzzles. These suggest a high-density emitting core around the WeSb 1 nucleus, making it a double-envelope nebula. Such a configuration bears a close resemblance to a rare class of PN, where the central core hosts a high-density compact emission knot (CEK). The formation channels of CEKs are not yet well understood. The leading hypothesis is the fall-back of some of the nebular ejecta on the white dwarf or its wide-orbit companion. The prototypical system is EGB 6 (Liebert et al., 2013; Bond et al., 2016), with only around a dozen more known of its kind (see Frew et al. 2016, and the recent discoveries by Bond et al. 2024b, a). The high electron densities (resulting in quenched [O III] lines and absence of [O II], [N II], and [S II] lines in spectra), near-IR excess, and the presence of a companion in WeSb 1 are all consistent with the properties of a EGB 6-type PN nucleus.
@@
We compare WeSb 1 with other systems known to show dust obscuration and find a similarity to NGC 2346. It is a binary system with one of the longest known periods inside a PN (16 days, see Mendez & Niemela 1981; Brown et al. 2019) and known to show dust obscuration (Mendez et al., 1982). The binary consists of a hot CSPN and an A-type subgiant (see Gómez-Muñoz et al. 2019) with stellar parameters comparable to those estimated for WeSb 1 in this work. Unlike WeSb 1, however, NGC 2346 does not show a high-density emission-line core, nor do they have the same nebular morphology (the latter is bipolar). We note here though that it is possible for WeSb 1 nebula to be bipolar, with the viewing angle being face-on. However, this poses an additional challenge of explaining a polar dust disk which can cause the transits. WeSb 1 also bears similarities to PN M 2-29 (Hajduk et al., 2008; Miszalski et al., 2011a) which show both dust obscuration and an emission-line core. In fact, the latter work discusses the same two possibilities for PN M 29 (EGB 6-type PN or D’-type symbiotic) as has been discussed in this work for WeSb 1. However, one difference remains: the transits in WeSb 1 are much shorter-lived (a few days) compared to the others (a few years).
@@ 5 Discussion and Conclusions
We have conducted a systematic study of the optical photometric variability of cataloged central stars of planetary nebulae (CSPNe, CW21) using the epochal photometric data from the Zwicky Transient Facility (ZTF). The main challenge in studying PNe with ground-based telescopes is atmospheric seeing. This often prevents adequate resolution of the CSPN from the surrounding nebula, introducing uncertainties (see Section 2.2). We employed stringent photometric quality cuts (Section 2.3) and used appropriate variability metrics (Normalized Excess Variance, see Section 2.4) to mitigate such effects. We arrived at a final list of highly variable CSPNe (the HNEV sample). The variables can broadly be classified into two qualitative classes: short-timescale and long-timescale, based on the temporal resolution of the variability with ZTF cadence. In this paper, we focus on the former class. The latter is the subject of Paper II (in prep.). The list of short-timescale variables is presented in Table H.
@@
Most (83) of the HNEV CSPNe belong to the class of short-timescale variables. We searched for periodicity using the Lomb-Scargle algorithm (Section 3.2). We recovered seventeen known periodic variables. Among them, the correct period was detected in twelve and half the true period in another three. We also find six new periodic sources, with confident period detection in four of them (also see Appendix B for three more sources among the non-HNEV objects). Among the aperiodic light curves, the majority are jittering sources showing irregular variability with flux amplitude in the range of 10 – 40%. Possible sources of such variability include wind from the hot CSPN, unresolved pulsations and/or local stochastic dust activity. We also find several outbursting sources. A literature search shows that we have recovered several objects with prior reports of photometric variability, including a significant fraction of the Gaia variables reported in Chornay et al. (2021).
@@
The variability in one of the objects, WeSb 1, appears unique. The optical light curves show very deep and irregular dips that resemble transits from debris disk (see Figure 4). The ingress and egress durations of the transits range over a few days (with some of them being well-resolved in ZTF cadence) with short-duration minima. Similar behavior is also witnessed in the near-IR -band light curve from Palomar Gattini-IR. Significant variability is also seen in Gaia and WISE light curves, though their cadence is inadequate in resolving the short-timescale activities. Such peculiarity prompted additional studies (see section 4). We find strong evidence for the binarity of the system. The optical spectrum and the SED constrain the spectral type of the companion roughly between A and G, with the best estimate being an F-type subgiant. With this knowledge, we use stellar evolutionary models to constrain the initial mass and the age of the system (see Section 4.4 and also Appendix F). The SED shows significant IR excess starting from near-IR wavelengths, indicative of a complex dust structure around the nucleus. Most of the line emissions appear confined to the nucleus, with the line ratios suggesting very high electron densities (). This is unexpected given the evolved morphology of the surrounding nebula. This makes WeSb 1 a candidate for either EGB 6-type PN hosting compact emission knots, or a symbiotic system inside an evolved PN, both of which are rare systems. A more detailed study on this object with additional data is in preparation.
@@
We conclude by briefly discussing the prospects of using ground-based photometric sky surveys to study CSPNe. With much higher cadence and longer temporal baselines compared to most sky-telescopes, these are ideal for studying the photometric variability of sources and resolving small timescale activities. This work (Paper I) and Paper II (in prep.), demonstrate the success of the existing facilities (primarily ZTF, but also other surveys like Asteroid Terrestrial-impact Last Alert System (ATLAS) in optical and Palomr Gattini-IR in near-IR) to identify unique sources. The upcoming Large Synoptic Survey Telescope (LSST, Ivezić et al. 2019) will widen the window of opportunity with its access to the galactic center where most of the PNe are situated. Effective variability metric(s) need to be identified to find ‘interesting’ sources from the pool of data (see Appendix A for such an example of a two-dimensional metric space). However, most algorithms currently in existence focus on point-source photometry. With the extended (and often bright) nebula, this is unsuitable for the study of CSPNe. On the other hand, simple aperture photometry of the whole nebula is also not desirable as i) we are interested in the CSPN, and ii) contamination from several field stars in the galactic plane. Thus, sufficient care must be taken in using the data and interpreting the results. In the long run, a robust photometry methodology with proper nebular subtraction is desirable.
Appendix A A unique metric space for unique variables
@@
where , are the standard deviation and point-to-point scatter metrics for the light curves. The latter is defined as the root-mean-square of the difference of successive data points (in this case, magnitude values) in a time series. SkewP is defined as:
@@
We calculate the metric values with only the ZTF forced photometry data. Similar to the band-average metric, we define the mean metric values from the and bands as and . Figure A1 shows the distribution of the objects in this metric space. The median and trimmed standard deviation values for the population are and respectively. Note that the median is broadly consistent with this metric’s theoretically expected random value. For , the median and standard deviation values are and respectively. These are consistent with the expectation that most objects do not show skewed light curves.
@@
Appendix B Future Prospects: New Periodic Variables among the non-HNEV objects
Through the application of strict photometric quality cuts and selection through variability metrics, we have missed several potentially interesting sources appearing in the broader ZTF dataset. To motivate the need for further studies, we present here (Figure B1) three periodic variables which do not appear in the HNEV sample: Pa 13 (PN G041.4-09.6), IsWe 2 (PN G107.7+07.8), and M 2-50 (PN G097.6-02.4). The current HASH catalog lists all three objects as “True” PNe. To the best of our knowledge, this is the first report of periodicity in the three sources. This readily shows the scope of improvement in the study of CSPNe with ZTF. With regards to periodicity, a more careful study revealing more such sources will be presented in Tam et al. (in preparation).

### v2

@@ 4.3 Spectral Energy Distribution
We queried the SED of WeSb 1 using the online VOSA SED tool (Bayo et al., 2008, Figure 10, blue squares) with a 3′′ cross-match radius with all the available catalogs. The flux outputs are presented in Table 4. VOSA failed to load the GALEX near-ultra-violet (NUV) flux, which was separately queried from the GALEX catalog212121https://galex.stsci.edu/GR6/?page=mastform and manually added to the table.
First, we analyze the SED independent of the inferences about the stellar parameters/spectral type in the previous section. We consider three cases of extinction: only interstellar (222222Obtained from the online query service: http://argonaut.skymaps.info/query, implementing the Green et al. (2019) three-dimensional extinction maps), a representative value for the Balmer decrement (), and the Na-D absorption (, as an example of an extreme case of extinction).232323We note here that the EW of the Na-D absorption in the Coelho spectrum is . Subtracting this from the observed EW leads to a revised extinction estimation of , consistent with all the other estimates. We still briefly consider this case just to demonstrate the effect of a large extinction, in case it is underestimated. The first three panels of Figure 10 show the reddening-corrected SEDs for these three extinction values. A strong IR excess, starting from the near-infrared wavelengths is evident, likely arising from circumstellar dust disk and/or shell. The stellar component dominates in the optical wavelengths.
@@ 4.4 Discussion
The position of the star in the Gaia color-magnitude diagram (CMD), however, is not consistent with a single white dwarf. With Gaia and , the object is significantly redder and more luminous than a white dwarf. All the previous analyses substantiate this, making it more likely to be a binary with the optical flux dominated by a cooler companion (A – G type star) of the hot CSPN (a chance alignment of a cool field star and the hot CSPN is unlikely as the probability of such an occurence is very low, see Appendix F). Thus, the observed photometric variability can arise from a transiting circumbinary debris or a circumstellar debris disk around the companion. The near-total eclipses, however, demands that the obscuring material be nearly as large as the star itself. With the inferred stellar radius being , it is more likely that the transits are from very thick dust/debris clouds rather than solid objects like asteroids. We note here that dust activity from the star itself might also have resulted in irregular variability (with RCrB stars being extreme cases242424Note that the WeSb 1 nucleus is unlikely to be an RCrB or similar hydrogen deficient star due to the presence of prominent Balmer/Paschen absorption lines). However, the best estimate of the spectral type of the companion, F-type subgiant, makes this scenario unlikely as such stars are not expected to take part in significant dust production.
@@
The inferred temperature and radius of the companion can be used to estimate its mass and the age of the system. We query stellar evolutionary tracks calculated using the Modules for Experiments in Stellar Astrophysics (MESA, Paxton et al. 2011, 2013, 2015) from the online portal of MESA Isochrones and Stellar Tracks (MIST, Dotter 2016; Choi et al. 2016)252525https://waps.cfa.harvard.edu/MIST/index.html. We visually compare the models with the stellar parameters calculated in the previous sections. The result is presented in Figure 11. We find that the inferred radius and temperature values lie between stars of masses 2.2 – 2.4 (left panel in the figure). With a generous allowance for error, we arrive at a mass estimation of . Thus, we get a corresponding age of the star of Gyr (middle panel in the figure). Note that at this evolutionary stage, the stellar radius (and temperature) evolves rapidly with time. Thus, the age estimation is not particularly sensitive to the exact estimations of the stellar parameters. At this mass and age, the model predicts , consistent with the previous estimates. The position of the star in the Gaia CMD is also consistent with those predicted by the models in the inferred mass range (with an applied extinction correction of , see right panel of the same figure). As the PN is still visible, this implies that the PN progenitor left the main sequence only recently. Thus, its initial mass needs to be very similar (slightly larger) to its subgiant companion. We thus obtain estimates of both the total initial mass of the binary and its age. We also find that the example stellar parameters of the hot CSPN used in Figure 10 are also broadly consistent with this evolutionary stage (see Appendix H). We note here, however, that there is a chance that the companion is not naturally in the subgiant phase (as indicated by the radius), but is just inflated from capture of the nebular ejecta. In that case, the mass and age estimation will no longer be valid.
@@
The compactness of the emission lines and the high inferred electron densities (see section 4.2) are additional puzzles. These suggest a high-density emitting core around the WeSb 1 nucleus, making it a double-envelope nebula. Such a configuration bears a close resemblance to a rare class of PN, where the central core hosts a high-density compact emission knot (CEK). The formation channels of CEKs are not yet well understood. The leading hypothesis is the fall-back of some of the nebular ejecta on the white dwarf or its wide-orbit companion. The prototypical system is EGB 6 (Liebert et al., 2013; Bond et al., 2016), with only around a dozen more known of its kind (see Frew et al. 2016, and the recent discoveries by Bond et al. 2024b, a). The high inferred electron densities, near-IR excess, and the presence of a companion in WeSb 1 are all consistent with the properties of a EGB 6-type PN nucleus.
@@
We compare WeSb 1 with other systems known to show dust obscuration and find a similarity to NGC 2346. It is a binary system with one of the longest known periods inside a PN (16 days, see Mendez & Niemela 1981; Brown et al. 2019) and known to show dust obscuration (Mendez et al., 1982). The binary consists of a hot CSPN and an A-type subgiant (see Gómez-Muñoz et al. 2019) with stellar parameters comparable to those estimated for WeSb 1 in this work. Unlike WeSb 1, however, NGC 2346 does not show a high-density emission-line core, nor do they have the same nebular morphology (the latter is bipolar). We note here though that it is possible for WeSb 1 nebula to be bipolar, with the viewing angle being face-on. However, this poses an additional challenge of explaining a polar dust disk which can cause the transits. WeSb 1 also bears similarities to PN M 2-29 (Hajduk et al., 2008; Miszalski et al., 2011a) which show both dust obscuration and an emission-line core. In fact, the latter work discusses the same two possibilities for PN M 2-29 (EGB 6-type PN or D’-type symbiotic) as has been discussed in this work for WeSb 1. However, one difference remains: the transits in WeSb 1 are much shorter-lived (a few days) compared to the others (a few years).
@@ 5 Discussion and Conclusions
We have conducted a systematic study of the optical photometric variability of cataloged central stars of planetary nebulae (CSPNe, CW21) using the light curves from the Zwicky Transient Facility (ZTF). The main challenge in studying PNe with ground-based telescopes is atmospheric seeing. This often prevents adequate resolution of the CSPN from the surrounding nebula, introducing uncertainties (see Section 2.2). We employed stringent photometric quality cuts (Section 2.3) and used appropriate variability metrics (Normalized Excess Variance, see Section 2.4) to mitigate such effects. We arrived at a final list of highly variable CSPN candidates (the HNEV sample). The variables can broadly be classified into two qualitative classes: short-timescale and long-timescale, based on the temporal resolution of the variability with ZTF cadence. In this paper, we focus on the former class. The latter is the subject of Paper II (in prep.). The list of short-timescale variables is presented in Table J.
@@
Most (83) of the HNEV CSPNe belong to the class of short-timescale variables. We searched for periodicity using the Lomb-Scargle algorithm (Section 3.2). We recovered seventeen known periodic variables. Among them, the correct period was detected in twelve and half the true period in another three. We also find six new periodic sources, with confident period detection in five of them (also see Appendix D for three more sources among the non-HNEV objects). Among the aperiodic light curves, the majority are jittering sources showing irregular variability with flux amplitude in the range of 10 – 40%. Possible sources of such variability include wind from the hot CSPN, unresolved pulsations and/or local stochastic dust activity. We also find several outbursting sources. A literature search shows that we have recovered several objects with prior reports of photometric variability, including a significant fraction of the Gaia variables reported in Chornay et al. (2021).
@@
The variability in one of the objects, WeSb 1, appears unique. The optical light curves show very deep and irregular dips that resemble transits from debris disk (see Figure 4). The ingress and egress durations of the transits range over a few days with short-duration minima. The transits are well-resolved in the available TESS light curves. Similar behavior is also witnessed in the near-IR -band light curve from Palomar Gattini-IR. Significant variability is also seen in Gaia and WISE light curves, though their cadence is inadequate in resolving the short-timescale activities. Such peculiarity prompted additional studies (see section 4). We find strong evidence for the binarity of the system. The optical spectrum and the SED constrain the spectral type of the companion roughly between A and G, with the best estimate being an F-type subgiant. With this knowledge, we use stellar evolutionary models to constrain the initial mass and the age of the system (see Section 4.4 and also Appendix H). The SED shows significant IR excess starting from near-IR wavelengths, indicative of a complex dust structure around the nucleus. Most of the line emissions appear confined to the nucleus, with the line ratios suggesting very high electron densities (). This is unexpected given the evolved morphology of the surrounding nebula. This makes WeSb 1 a candidate for either EGB 6-type PN hosting compact emission knots, or a symbiotic system inside an evolved PN, both of which are rare systems. A more detailed study on this object with additional data is in preparation.
@@
We conclude by briefly discussing the prospects of using ground-based photometric sky surveys to study CSPNe. With much higher cadence and longer temporal baselines compared to most sky-telescopes, these are ideal for studying the photometric variability of sources and resolving small timescale activities. This work (Paper I) and Paper II (in prep.), demonstrate the success of the existing facilities (primarily ZTF, but also other surveys like Asteroid Terrestrial-impact Last Alert System (ATLAS) in optical and Palomr Gattini-IR in near-IR) to identify unique sources. The upcoming Large Synoptic Survey Telescope (LSST, Ivezić et al. 2019) will widen the window of opportunity with its access to the galactic center where most of the PNe are situated. Effective variability metric(s) need to be identified to find ‘interesting’ sources from the pool of data (see Appendix B for such an example of a two-dimensional metric space). However, most algorithms currently in existence focus on point-source photometry. With the extended (and often bright) nebula, this is unsuitable for the study of CSPNe. On the other hand, simple aperture photometry of the whole nebula is also not desirable as i) we are interested in the CSPN, and ii) contamination from several field stars in the galactic plane. Thus, sufficient care must be taken in using the data and interpreting the results. In the long run, a robust photometry methodology with proper nebular subtraction is desirable.
Appendix A ZTF Data and the NEV Metric
The inclusion of this section is largely motivated from the comments of the anonymous reviewer. Here, we continue our discussion from Section 2.4 and further elaborate on the NEV metric and its application to the ZTF data. First, we provide the statistics on the number of data points per object, on which we apply our variability metric. The top panel of Figure A1 shows the distribution of the number of data points for each object which passes all the photometric quality cuts for both the forced photometry (FP) and standard photometry (SP). For both bands, the median number of data points per object is , after rejecting all the object with less than “good” data points. This shows that, for most objects, we have sufficient number of data points for a reliable evaluation of the variability metric.
The distribution of the number of data points, however, is not narrow. This leads to the question of how much the metric NEV gets artificially affected by the number of data points. In the middle panel of the same figure, we plot the NEV metrics for all the objects as a function of their number of good data points, . The scatter plot shows no correlation. The NEV metric is indeed a normalized metric and does not depend on (the authors have also verified this through synthetic data). To highlight any trend, we compute the median of the NEVs within logarithmic bins of width dex in . A weak positive correlation is seen with the FP data. However, we reason that this is a “true” trend and not an artifact. Objects with smaller number of data points are also likely the objects with high uncertainty in photometry (which led to rejection of several of the data points). Thus, they will tend to have lower NEV metric values. With the SP data, however, we do not see the trend. The medians, however, are skewed to higher values. This is due to the underestimation of errors for SP as discussed in Section 2.4.
We note here that, though NEV is not dependent on , it is indeed strongly dependent on the photometric errors. This is one of main reasons why this metric has been chosen, given the often large errors seen in photometry of PNe (see discussion in Section 2.2). But this may result in false positives through the uncertainties in the photometric errors themselves. This is prevented through the use of both standard and forced photometry data, which undergo different processing steps and independent error assignments. Thus, an agreement between the two is a strong indication of the true variability of the object.
In the last panel of Figure A1 we justify our choice of NEV over other commonly used metrics like standard deviation (SD). The left (right) panel shows SD (NEV) as a function of the Gaia magnitude of the objects. We only use FP data for an even comparison. We clearly see that SD suffers from a increasing trend with magnitude, owing to larger photon noise for fainter objects. The trend disappears with NEV. On the contrary, the fainter objects are seen to have lower NEV values. This is also a true trend, as fainter objects are expected to have poorer photometry (and thus larger errors, and lower NEV). This discussion justifies the suitability of NEV as a variability metric for this work.
Appendix B A unique metric space for unique variables
@@
where and are the standard deviation and point-to-point scatter metrics for the light curves, respectively. The latter is defined as the root-mean-square of the difference of successive data points (in this case, magnitude values) in a time series. SkewP is defined as:
@@
We calculate the metric values with only the forced photometry data. Similar to the band-average metric, we define the mean metric values from the and bands as and . Figure B1 shows the distribution of the objects in this metric space. The median and trimmed standard deviation values for the population are and respectively. Note that the median is broadly consistent with this metric’s theoretically expected random value. For , the median and standard deviation values are and respectively. These are consistent with the expectation that most objects do not show skewed light curves.
@@
Appendix C Follow-up Observations with the Liverpool Telescope
The Liverpool Telescope observations were obtained through a Sloan filter with the IO:O instrument in queue mode under program ID XPL21A08 between October 2021 and February 2022. The data were reduced by the standard IO:O pipeline and then differential photometry of the central stars was performed against non-variable field stars using standard astropy routines (Astropy Collaboration et al., 2018)
We performed an independent Lomb-Scargle period search on the LT data over the same period range and with the same resolution as mentioned in Section 3.2. We obtain periods well consistent with those inferred from the ZTF data. This serves as a confirmation for the periods, especially for Kn 51 and HaWe 8, where the ZTF data is noisy and the period is close to 1 day (ZTF cadence), respectively. The phase-folded data are presented in Figure C1. For even comparison with the ZTF data, the same epoch has been chosen as the starting phase. The apparent phase shift of the LT data of Kn 51 with respect to the ZTF data, when phase-folded at the LT-inferred period, is an artifact, owing to the small difference (2 seconds) in the inferred periods from the two light curves. The difference is unlikely to be real or significant.
Appendix D Future Prospects: New Periodic Variables among the non-HNEV objects
Through the application of strict photometric quality cuts and selection through variability metrics, we have missed several potentially interesting sources appearing in the broader ZTF dataset. To motivate the need for further studies, we present here (Figure D1) three new periodic variables which do not appear in the HNEV sample: Pa 13 (PN G041.4-09.6), IsWe 2 (PN G107.7+07.8), and M 2-50 (PN G097.6-02.4). The current HASH catalog lists all three objects as “True” PNe. To the best of our knowledge, this is the first report of periodicity in these three sources. This readily shows the scope of improvement in the study of CSPNe with ZTF. With regards to periodicity, a more careful study revealing more such sources will be presented in Tam et al. (in preparation).